OpenStudy (anonymous):

What are the solutions of 6x2 − x − 2 = 0? x = 1/2, x = -2/3 x = -1/2, x = 2/3 x = 1/6, x = −1 x = −1/6, x = 2

OpenStudy (anonymous):

Try using the quadratic formula: For a quadratic in the form: \[ax ^{2}+bx+c=0\] \[x=\frac{ -b \pm \sqrt{b ^{2}-4*a*c} }{ 2*a }\] Hope that helps!

OpenStudy (anonymous):

I think so, I got x= -1/2 , x= 2/3 .... is that correct?

OpenStudy (anonymous):

Yes, correct! Good work :)

OpenStudy (anonymous):

Thank You so much :)

OpenStudy (anonymous):

You're very welcome